# Analysis of Decentralized Stochastic Successive Convex Approximation for composite non-convex problems

### v1

@@ II-D Approximate optimality
where $\hat{{\mathbf{w}}}_{i}^{t}\in\partial(h+11_{{\mathcal{X}}})({\hat{{\mathbf{x}}}}_{i}^{t})$. If (6) holds then the output $\tilde{{\mathbf{x}}}$ of D-MSSCA is chosen uniformly at random from the set $\{{\hat{{\mathbf{x}}}}_{i}^{t}\}_{0\leq t\leq T}^{i\in{\mathcal{V}}}$ then we have ${\mathbb{E}}\left[\left\|\nabla u(\tilde{{\mathbf{x}}})+\hat{{\mathbf{w}}}_{\tilde{{\mathbf{x}}}}\right\|^{2}\right]\leq\epsilon$.
@@ III Convergence Analysis
In this section, we will provide a detailed convergence analysis of the D-MSSCA algorithm and compare its rate with other state-of-the-art algorithms. For the analysis, we have defined the concatenated vectors ${\mathbf{x}},{\mathbf{y}},{\mathbf{z}},{\hat{{\mathbf{x}}}}\in{\mathbb{R}}^{nd}$ by concatenating corresponding local vectors $\{{\mathbf{x}}_{i}\in{\mathbb{R}}^{d}\}_{i\in{\mathcal{V}}},\{{\mathbf{y}}_{i}\in{\mathbb{R}}^{d}\}_{i\in{\mathcal{V}}},\{{\mathbf{z}}_{i}\in{\mathbb{R}}^{d}\}_{i\in{\mathcal{V}}},\{{\hat{{\mathbf{x}}}}_{i}\in{\mathbb{R}}^{d}\}_{i\in{\mathcal{V}}}$ of all the nodes. Utilizing these concatenated vectors, we can write the update equation (3) and (5) in a more compact form as below,
@@
where, ${\underline{{\mathbf{W}}}}={\mathbf{W}}\otimes{\mathbf{I}}_{d}\in{\mathbb{R}}^{nd\times nd}.$ Furthermore, we have defined the concatenated local gradient vector $\nabla u({\mathbf{x}}^{t}):=[\nabla u_{1}({\mathbf{x}}_{1}^{t})^{\mathsf{T}},\nabla u_{2}({\mathbf{x}}_{1}^{t})^{\mathsf{T}},\cdots,\nabla u_{n}({\mathbf{x}}_{n}^{t})^{\mathsf{T}}]^{\mathsf{T}}\in{\mathbb{R}}^{nd}$, where $\nabla u_{i}({\mathbf{x}}_{i}^{t})\in{\mathbb{R}}^{d}$ for all $i\in{\mathcal{V}}$. For the sake of brevity, we define for all $t$
@@
We began our analysis by stating some standard results used in decentralized optimization in Lemma 1 [26, 5, 31]. Next, Lemmm 2, Lemma 3 and Lemma 4 establish the contraction relations for $\theta^{t},\phi^{t},\upsilon^{t}$ and $\varepsilon^{t}$. Proceeding further, using these contraction results, we bounded the cumulative error accumulation $\sum_{t=1}^{T}{\mathbb{E}}\left[\theta^{t}\right],\sum_{t=1}^{T}\phi^{t},\sum_{t=1}^{T}\varepsilon^{t}$ and $\sum_{t=1}^{T}\upsilon^{t}$ in Lemma 6, Lemma 7, and Lemma 8. Lemma 7, then extends these findings to provide bounds on the accumulated average iterate progress $\sum_{t=1}^{T}{\mathbb{E}}\left[\delta^{t}\right]$ under certain step-size conditions. Finally, Theorem 1 uses all these results to quantify the SFO complexity of D-MSSCA.
The results presented in Lemma 2-8 are can be obtained by applying the findings of [26] to SCA. However, our bounds are different as we have defined the quantities in terms of ${\boldsymbol{\delta}}$ rather than $\upsilon$, and the variations in the intermediate steps are detailed in the proof. The results in Lemma 9 and Lemma 10 are different from [26] due to our focus on average progress $\Delta^{T}=\frac{1}{T}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}^{t}\right\|^{2}\right]$. Results similar to those in Lemma 10 are presented in [8]; however, in that work, the bounds are achieved for a deterministic case. The results of Lemma 10 and Theorem 1 are entirely novel to the considered class of problem.
@@
Under Assumptions  A4 and A7, we have the following results for all $t\geq 1$, where ${\mathbf{x}}^{t},{\mathbf{y}}^{t},{\mathbf{z}}^{t}$ are variables of D-MSSCA at iterate $t$.
@@
where ${\lambda_{{\mathbf{W}}}}:=\lambda_{\max}({\mathbf{W}}-\frac{1}{n}\mathbf{1}\mathbf{1}^{\mathsf{T}})$. The proofs of the above results are straightforward and can be found in [12, 32]. In the proof of (14a), the contraction property of doubly stochastic symmetric matrices is used; in (14b), the smoothness of the functions $u_{i}$ ($i\in{\mathcal{V}}$); in (14c), the special initialization condition of $v^{t}$ and the doubly stochastic property of ${\mathbf{W}}$; and in (14d), the property of norm with the Cauchy-Schwarz inequality. The next Lemma bounds the consensus errors in the ${\mathbf{x}}^{t}-$updates (7) of the D-MSSCA algorithm.
@@
Under Assumption (A7), for the ${\mathbf{x}}^{t}-$ updates of D-MSSCA algorithm, the following inequality holds for all $t\geq 2$ and $\eta_{1}>0$
@@
The proof of Lemma 2 is provided in Appendix A and follows by applying update step (7), then separating the terms using Young’s inequality. Finally, by utilizing the properties of the communication matrix ${\mathbf{W}}$ (14a); we obtain the desired results. Similar contraction bounds on $\theta^{t}$ have been achieved in various gradient-tracking based decentralized optimization algorithms [32, 26]. However, our bound is slightly different because we define it in terms of $\theta^{t-1}$ and $\left\|{\boldsymbol{\delta}}^{t-1}\right\|^{2}$ instead of $\theta^{t-1}$ and $\upsilon^{t-1}$, as seen in literature. Next, we will bound the gradient variances.
@@
Under Assumptions A2-A4, and Assumption A7, The following inequalities hold for the iterates produced by D-MSSCA algorithm, where $t\geq 2,0<\alpha<1,\eta_{1},\eta_{2},\eta_{3}>0$
@@
$\upsilon^{t}$ $\leq(1-\beta)^{2}\upsilon^{t-1}+3L^{2}(1-\beta)^{2}\left(1+\frac{1}{\eta_{3}}\right)\left[\left(\theta^{t}\right)^{2}+n\alpha^{2}\left\|\left(\frac{1}{n}\mathbf{1}^{\mathsf{T}}\otimes{\mathbf{I}}_{d}\right){\boldsymbol{\delta}}^{t-1}\right\|^{2}+\left(\theta^{t-1}\right)^{2}\right]$
@@
The proof of Lemma 3 proceeds along the similar lines as the proof in [26, Lemma 3]. However, our final bounds are in terms of $\bar{{\boldsymbol{\delta}}}^{t-1}$ rather than ${\bar{{\mathbf{z}}}}^{t-1}$ in [26]. This difference is due to the ${\mathbf{x}}-$update of D-MSSCA (3), which differs from GT-HSGD algorithm proposed in [26]. The proof uses the unbiased nature of the local gradient estimate ${\mathbf{z}}_{i}^{t}$. Furthermore, as the gradient estimate at each node is independent of those at other nodes given the history sequence ${\mathcal{H}}^{t}$, we can omit the cross terms of inner products appearing in the intermediate steps to obtain simplified expressions. Finally, by applying Assumption A4 alongside ${\mathbf{x}}_{i}^{t}$ updates, we get the desired result.
@@
Under Assumptions A2-A4 and A7, the following inequality holds for $\beta\in(0,1)$, $\forall t\geq 2$,
@@
The above lemma bounds the error in local estimation of the global gradient. The proof uses the ${\mathbf{y}}^{t}-$update (8), applies conditional expectation, and simplifies intermediate steps using Assumptions A2,A3, and A4. Finally, by utilizing the consensus error bound in Lemma 2 and few mathematical simplification we achieved the desired result. The elaborated proof can be found in Appendix B. The next Lemma provides the basic results of non-negative sequences, which are necessary to bound the error accumulation in subsequent lemmas.
@@
    $\sum_{t=1}^{T}V^{t}\leq\frac{V_{1}}{1-q}+\frac{\sum_{t=2}^{T}Q^{t-1}}{1-q}+\frac{CT}{1-q}.$ (18)
The above mentioned recursion results align with [26, Lemma 6], and the proof follows a similar approach. For the complete proof, refer to [26]. Using the stated lemmas, we can now establish upper bounds for the cumulative errors up to iteration $T$, as detailed in the following lemmas.
@@
The proof of Lemma 9 begins by defining the optimality condition of (1), and utilizing it with the properties of surrogate to get the descent direction. Further by using the convexity of ${\mathbf{h}}$ and the properties of the communication matrix ${\mathbf{W}}$, along with some mathematical simplifications, we achieve the desired result. The complete proof of Lemma 9 is detailed in Appendix F.
@@
On combining the common terms and substituting the bound of $\sum_{t=1}^{T}\varepsilon^{t}$ from Lemma 8, we get
@@
On defining $C_{\mu}=\Bigg{[}\frac{6\alpha^{2}{\lambda_{{\mathbf{W}}}}^{2}\gamma_{1}L^{2}}{\left(1-{\lambda_{{\mathbf{W}}}}^{2}\right)^{2}}\left(19+12\beta+\frac{12}{\beta n}+\frac{24{\lambda_{{\mathbf{W}}}}^{2}}{\left(1-{\lambda_{{\mathbf{W}}}}^{2}\right)^{2}}\left(3+4\beta\right)\right)\mu+\frac{1}{2\gamma_{1}}+\frac{\alpha L}{2}+\frac{9L^{2}\alpha^{2}\gamma_{1}}{n\beta}\Bigg{]}$, we have:
@@
where in (i), we used the fact that ${\boldsymbol{\xi}}_{i}^{1,l},{\boldsymbol{\xi}}_{j}^{1,m}$ are independent for $l\neq m$ and in (ii), we applied (A3). On substituting the bound $\varepsilon^{1}\leq\frac{2\bar{\sigma}^{2}}{b_{0}^{2}}+2\left\|\nabla u({\mathbf{x}}^{1})\right\|^{2},$ and $U^{\star}<U({\bar{{\mathbf{x}}}}^{T+1})$, we obtain:,
@@
Further if we consider $\beta=\alpha^{2}$,$\gamma_{1}=\frac{1}{\mu}$, and $0<\alpha\leq\min\Bigg{\{}\frac{1}{114},\frac{\left(1-{\lambda_{{\mathbf{W}}}}^{2}\right)^{2}}{432{\lambda_{{\mathbf{W}}}}^{2}},\left(\frac{\left(1-{\lambda_{{\mathbf{W}}}}^{2}\right)}{24{\lambda_{{\mathbf{W}}}}}\right)^{2/3},\frac{\mu}{6L},\frac{\mu^{2}\left(1-{\lambda_{{\mathbf{W}}}}^{2}\right)^{2}}{48L^{2}{\lambda_{{\mathbf{W}}}}^{2}}\Bigg{\}}$, and if $\mu\geq\frac{6\sqrt{3}L}{n}\left(1+\frac{8{\lambda_{{\mathbf{W}}}}^{2}}{(1-{\lambda_{{\mathbf{W}}}}^{2})}\right)$, then after further simplification we get $C_{\mu}\geq\frac{\mu}{4}>0$, utilizing which we get the desired result. ∎
@@
On adding and subtracting $\nabla{\hat{f}}_{i}\left({\mathbf{x}}_{i}^{t},{\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)$, applying the definition of $\pi_{i}^{t}$ and update ${\mathbf{z}}_{i}^{t}$ (4), we obtain
@@
In (i), we have utilized assumption A4, and in (ii), applied (14b) and (14c). Now, by substituting the value of $\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{\bar{\nabla u}}({\mathbf{x}}^{t})-{\bar{{\mathbf{z}}}}^{t}\right\|^{2}\right]$, from Lemma 7, we get
@@
Further on combining the common terms and substituting the bound of $\varepsilon^{t}$ from Lemma 8, we get
@@
On combining the common terms and substituting the bound of $\sum_{t=1}^{T}{\mathbb{E}}\left[\left(\theta^{t}\right)^{2}\right]$ from Lemma 6, we can further simplify as follows:
@@
Further on substituting $\beta=\alpha^{2}$, considering $\alpha\leq\min\Bigg{\{}\frac{1}{116},\frac{\left(1-{\lambda_{{\mathbf{W}}}}^{2}\right)^{2}}{432{\lambda_{{\mathbf{W}}}}^{2}},\frac{\left(1-{\lambda_{{\mathbf{W}}}}^{2}\right)^{2/3}}{8{\lambda_{{\mathbf{W}}}}^{2/3}},\frac{\left(1-{\lambda_{{\mathbf{W}}}}^{2}\right)^{2}}{4{\lambda_{{\mathbf{W}}}}^{2}}\Bigg{\}}$, and rearranging, we get the desired result. ∎
@@
The SFO complexity of Corollary 1 matches that of DEEPSTORM [7] and ProxGT-SR-O/E [5]. It should be noted that, unlike DEEPSTORM and ProxGT-SR-O/E which require small and large batch sizes respectively, our algorithm is batchless and uses one sample at each iteration. Also, this rate matches the SFO complexity lower bound for centralized unconstrained stochastic non-convex optimization problems.

### v2

@@ II-D Approximate optimality
where $\hat{{\mathbf{w}}}_{i}^{t}\in\partial(h+11_{{\mathcal{X}}})\mid_{{\mathbf{x}}={\hat{{\mathbf{x}}}}_{i}^{t}}$. If (6) holds then the output $\tilde{{\mathbf{x}}}$ of D-MSSCA is chosen uniformly at random from the set $\{{\hat{{\mathbf{x}}}}_{i}^{t}\}_{0\leq t\leq T}^{i\in{\mathcal{V}}}$ then we have ${\mathbb{E}}[\|\nabla u(\tilde{{\mathbf{x}}})+\hat{{\mathbf{w}}}_{\tilde{{\mathbf{x}}}}\|^{2}]\leq\epsilon$ where $\hat{{\mathbf{w}}}_{\tilde{{\mathbf{x}}}}\in\partial(h+11_{{\mathcal{X}}})\mid_{{\mathbf{x}}=\tilde{{\mathbf{x}}}}$.
@@ III Convergence Analysis
This section will provide a detailed convergence analysis of the D-MSSCA algorithm and compare its rate with other state-of-the-art algorithms. For the analysis, we define the concatenated vectors ${\mathbf{x}},{\mathbf{y}},{\mathbf{z}},{\hat{{\mathbf{x}}}}\in{\mathbb{R}}^{nd}$ by concatenating corresponding local vectors $\{{\mathbf{x}}_{i}\in{\mathbb{R}}^{d}\}_{i\in{\mathcal{V}}},\{{\mathbf{y}}_{i}\in{\mathbb{R}}^{d}\}_{i\in{\mathcal{V}}},\{{\mathbf{z}}_{i}\in{\mathbb{R}}^{d}\}_{i\in{\mathcal{V}}},\{{\hat{{\mathbf{x}}}}_{i}\in{\mathbb{R}}^{d}\}_{i\in{\mathcal{V}}}$ of all the nodes. Using these concatenated vectors, we can write the update equation (3) and (5) in a more compact form as below,
@@
where, ${\underline{{\mathbf{W}}}}={\mathbf{W}}\otimes{\mathbf{I}}_{d}\in{\mathbb{R}}^{nd\times nd}.$ Furthermore, we define the concatenated local gradient vector $\nabla u({\mathbf{x}}^{t}):=[\nabla u_{1}({\mathbf{x}}_{1}^{t})^{\mathsf{T}},\nabla u_{2}({\mathbf{x}}_{1}^{t})^{\mathsf{T}},\cdots,\nabla u_{n}({\mathbf{x}}_{n}^{t})^{\mathsf{T}}]^{\mathsf{T}}\in{\mathbb{R}}^{nd}$, where $\nabla u_{i}({\mathbf{x}}_{i}^{t})\in{\mathbb{R}}^{d}$ for all $i\in{\mathcal{V}}$. For the sake of brevity, we define for all $t$
@@
We begin our analysis by stating some standard results used in decentralized optimization in Lemma 1 [25, 5, 32]. Next, Lemmm 2, Lemma 3 and Lemma 4 establish the contraction relationships for $\theta^{t},\phi^{t},\upsilon^{t}$ and $\varepsilon^{t}$. These contraction relationships help in bounding the cumulative error accumulation $\sum_{t=1}^{T}{\mathbb{E}}\left[\theta^{t}\right],\sum_{t=1}^{T}\phi^{t},\sum_{t=1}^{T}\varepsilon^{t}$ and $\sum_{t=1}^{T}\upsilon^{t}$ in Lemma 6, Lemma 7, and Lemma 8. Lemma 7, then extends these findings to obtain bounds on the accumulated average iterate progress $\sum_{t=1}^{T}{\mathbb{E}}\left[\delta^{t}\right]$ under certain step-size conditions. Finally, Theorem 1 uses all these results to quantify the SFO complexity of D-MSSCA.
The results presented in Lemma 2-8 can be obtained by applying the approach of [25] to the SCA framework. However, our bounds are different as we define the quantities in terms of ${\boldsymbol{\delta}}$ rather than $\upsilon$, and the variations in the intermediate steps are detailed in the proof. The results in Lemma 9 and Lemma 10 are different from [25] due to our focus on average progress $\Delta^{T}=\frac{1}{T}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}^{t}\right\|^{2}\right]$. The results of Lemma 10 and Theorem 1 are entirely novel to this work. It is worth noting that the bound in Lemma 10 is similar to the deterministic bound in [8]. However, the steps needed to obtain those bounds have been omitted in [8].
@@
Under Assumptions A4 and A7, we have the following results for all $t\geq 1$, where ${\mathbf{x}}^{t},{\mathbf{y}}^{t},{\mathbf{z}}^{t}$ are variables of D-MSSCA at iterate $t$.
@@
where ${\lambda_{{\mathbf{W}}}}:=\lambda_{\max}({\mathbf{W}}-\frac{1}{n}\mathbf{1}\mathbf{1}^{\mathsf{T}})$. The proofs of the above results are straightforward and can be found in [12, 30]. In the proof of (14a), the contraction property of doubly stochastic symmetric matrices is used. In (14b), the smoothness of the functions $u_{i}$ ($i\in{\mathcal{V}}$) is applied. In (14c), the special initialization condition of $v^{t}$ and the doubly stochastic property of ${\mathbf{W}}$ are used. Lastly, in (14d), the norm property along with the Cauchy-Schwarz inequality is applied. The next Lemma bounds the consensus errors in the ${\mathbf{x}}^{t}-$updates (7) of the D-MSSCA algorithm.
@@
Under (A7), for the ${\mathbf{x}}^{t}-$ updates of D-MSSCA algorithm, the following inequality holds for all $t\geq 2$ and $\eta_{1}>0$
@@
The proof of Lemma 2 is provided in Appendix A and follows by applying update step (7), then separating the terms using Young’s inequality. Finally, by using the properties of the communication matrix ${\mathbf{W}}$ (14a); we obtain the desired results. Similar contraction bounds on $\theta^{t}$ have been achieved in various gradient-tracking-based decentralized optimization algorithms [30, 25]. However, our bound is slightly different because we define it in terms of $\theta^{t-1}$ and $\left\|{\boldsymbol{\delta}}^{t-1}\right\|^{2}$ instead of $\theta^{t-1}$ and $\upsilon^{t-1}$, as seen in the literature. Next, we will bound the gradient variances.
@@
Under A2-A4, and A7, The following inequalities hold for the iterates produced by D-MSSCA algorithm, where $t\geq 2,0<\alpha<1,\eta_{1},\eta_{2},\eta_{3}>0$
@@
$\upsilon^{t}$ $\leq(1-\beta)^{2}\upsilon^{t-1}+3L^{2}(1-\beta)^{2}\left(1+\frac{1}{\eta_{3}}\right){\mathbb{E}}\left[\left(\theta^{t}\right)^{2}+n\alpha^{2}\left\|\left(\frac{1}{n}\mathbf{1}^{\mathsf{T}}\otimes{\mathbf{I}}_{d}\right){\boldsymbol{\delta}}^{t-1}\right\|^{2}+\left(\theta^{t-1}\right)^{2}\right]$
@@
The proof of Lemma 3 is identical to that of [25, Lemma 3]. However, our final bounds are in terms of $\bar{{\boldsymbol{\delta}}}^{t-1}$ rather than ${\bar{{\mathbf{z}}}}^{t-1}$ in [25]. This difference is due to the ${\mathbf{x}}-$update of D-MSSCA (3), which differs from GT-HSGD algorithm proposed in [25]. The proof uses the unbiased nature of the local gradient estimate ${\mathbf{z}}_{i}^{t}$. Furthermore, as the gradient estimate at each node is independent of those at other nodes given the history sequence ${\mathcal{H}}^{t}$, we can omit the cross terms of inner products appearing in the intermediate steps to obtain simplified expressions. Finally, by applying A4 alongside ${\mathbf{x}}_{i}^{t}$ updates, we get the desired result.
@@
Under A2-A4 and A7, the following inequality holds for $\beta\in(0,1)$, $\forall t\geq 2$,
@@
The above lemma bounds the error in local estimation of the global gradient. The proof uses the ${\mathbf{y}}^{t}-$update (8), applies conditional expectation, and simplifies intermediate steps using Assumptions A2,A3, and A4. Finally, by using the consensus error bound in Lemma 2 and performing a few mathematical simplifications, the desired result is achieved. The elaborated proof can be found in Appendix B. The next Lemma provides the basic results of non-negative sequences, which are necessary to bound the error accumulation in subsequent lemmas.
@@
    $\sum_{t=1}^{T}V^{t}\leq\frac{V^{1}}{1-q}+\frac{\sum_{t=2}^{T}Q^{t-1}}{1-q}+\frac{CT}{1-q}.$ (18)
The above mentioned recursion results align with [25, Lemma 6], and the proof follows a similar approach. However, there is a slight difference between our results and theirs, due to the variation in the relationship of sequences and the range of $T$ considered here. For the complete proof, refer to [25]. Using the stated lemmas, upper bounds for the cumulative errors up to iteration $T$ can now be established, as detailed in the following lemmas.
@@
The proof of Lemma 9 begins by defining the optimality condition of (1), and using it with the properties of surrogate to get the descent direction. Further by using the convexity of ${\mathbf{h}}$ and the properties of the communication matrix ${\mathbf{W}}$, along with some mathematical simplifications, we achieve the desired result. The complete proof of Lemma 9 is detailed in Appendix F.
@@
Combining the common terms and substituting the bound of $\sum_{t=1}^{T}\varepsilon^{t}$ from Lemma 8, we get
@@
Defining $C_{\mu}=-\Bigg{[}\frac{6\alpha^{2}{\lambda_{{\mathbf{W}}}}^{2}\gamma_{1}L^{2}}{\left(1-{\lambda_{{\mathbf{W}}}}^{2}\right)^{2}}\left(19+12\beta+\frac{12}{\beta n}+\frac{24{\lambda_{{\mathbf{W}}}}^{2}}{\left(1-{\lambda_{{\mathbf{W}}}}^{2}\right)^{2}}\left(3+4\beta\right)\right)-\mu+\frac{1}{2\gamma_{1}}+\frac{\alpha L}{2}+\frac{9L^{2}\alpha^{2}\gamma_{1}}{n\beta}\Bigg{]}$, we have:
@@
where in (i), we used the fact that ${\boldsymbol{\xi}}_{i}^{1,l},{\boldsymbol{\xi}}_{j}^{1,m}$ are independent for $l\neq m$ and in (ii), we applied (A3). Substituting the bound $\varepsilon^{1}\leq\frac{2\bar{\sigma}^{2}}{b_{0}^{2}}+2\left\|\nabla u({\mathbf{x}}^{1})\right\|^{2},$ and $U^{\star}<U({\bar{{\mathbf{x}}}}^{T+1})$, we obtain:,
@@
Further if we consider $\beta=\alpha^{2}$,$\gamma_{1}=\frac{1}{\mu}$, and $0<\alpha\leq\min\Bigg{\{}\frac{1}{114},\frac{\left(1-{\lambda_{{\mathbf{W}}}}^{2}\right)^{2}}{432{\lambda_{{\mathbf{W}}}}^{2}},\left(\frac{\left(1-{\lambda_{{\mathbf{W}}}}^{2}\right)}{24{\lambda_{{\mathbf{W}}}}}\right)^{2/3},\frac{\mu}{6L},\frac{\mu^{2}\left(1-{\lambda_{{\mathbf{W}}}}^{2}\right)^{2}}{48L^{2}{\lambda_{{\mathbf{W}}}}^{2}}\Bigg{\}}$, and if $\mu\geq\frac{6\sqrt{3}L}{n}\left(1+\frac{8{\lambda_{{\mathbf{W}}}}^{2}}{(1-{\lambda_{{\mathbf{W}}}}^{2})}\right)$, then after further simplification we get $C_{\mu}\geq\frac{\mu}{4}>0$, using which we get the desired result. ∎
@@
Further adding and subtracting $\nabla{\hat{f}}_{i}\left({\mathbf{x}}_{i}^{t},{\mathbf{x}}_{i}^{t},\xi_{i}^{t}\right)$, applying the definition of $\pi_{i}^{t}$ and update ${\mathbf{z}}_{i}^{t}$ (4), we obtain
@@
In (i), we have applied assumption A4, and in (ii), we have applied (14b) and (14c). Now, by substituting the value of $\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{\bar{\nabla u}}({\mathbf{x}}^{t})-{\bar{{\mathbf{z}}}}^{t}\right\|^{2}\right]$, from Lemma 7, we get
@@
Further combining the common terms and substituting the bound of $\varepsilon^{t}$ from Lemma 8, we get
@@
Combining the common terms and substituting the bound of $\sum_{t=1}^{T}{\mathbb{E}}\left[\left(\theta^{t}\right)^{2}\right]$ from Lemma 6, we can further simplify as follows:
@@
Further substituting $\beta=\alpha^{2}$, considering $\alpha\leq\min\Bigg{\{}\frac{1}{116},\frac{\left(1-{\lambda_{{\mathbf{W}}}}^{2}\right)^{2}}{432{\lambda_{{\mathbf{W}}}}^{2}},\frac{\left(1-{\lambda_{{\mathbf{W}}}}^{2}\right)^{2/3}}{8{\lambda_{{\mathbf{W}}}}^{2/3}},\frac{\left(1-{\lambda_{{\mathbf{W}}}}^{2}\right)^{2}}{4{\lambda_{{\mathbf{W}}}}^{2}}\Bigg{\}}$, and rearranging, we get the desired result. ∎
@@
The SFO complexity of Corollary 1 matches that of DEEPSTORM [7] and ProxGT-SR-O/E [5]. It should be noted that, unlike ProxGT-SR-O/E, which requires large batch sizes, our algorithm is batchless and uses one sample at each iteration. Also, this rate matches the SFO complexity lower bound for centralized unconstrained stochastic non-convex optimization problems. Also, this rate matches the SFO complexity lower bound for centralized unconstrained stochastic non-convex optimization problems.